# Quantum-limited superresolution of two incoherent point sources with unknown photon numbers

### v1

@@ Abstract
Superresolution has been demonstrated to overcome the limitation of the Rayleigh’s criterion and achieve significant improvement of the precision in resolving the centroid separation of two incoherent optical point sources. However, in recent years, it was found that if the photon numbers of the two incoherent optical sources are unknown, the precision of superresolution vanishes when the two photon numbers are actually different. In this work, we analyze the estimation precision of the centroid separation between two incoherent optical sources with the same point-spread functions and unknown photon numbers in detail, and show that when the photon numbers of the two optical sources are different but sufficiently close, the superresolution can still realized but with different precisions. We find the condition on how close the photons numbers of two optical sources need to be to realize the superresolution, and derive the precision of superresolution in different regimes of the photon number difference. We further consider the superresolution for two incoherent optical sources with different point-spread functions, and show that the competition between the difference of photon numbers, the difference of the two point-spread functions and the centroid separation of the two optical sources determines the precision of superresolution. The results exhibit various precision limits of the centroid separation distinct from that of two point sources with identical point-spread functions and equal photon numbers. The results are finally illustrated by Gaussian point-spread functions.
@@ I Introduction
Optical imaging is an important sensing technique that has contributed to various research areas, such as medical technology, astronomy, nanotechnology and biochemistry, etc. Typical imaging methods include classical and quantum imaging [1, 2], fluorescence imaging [3, 4, 5], the superlens [6, 7], two-photon microscopy [8, 9], quantum holography [10, 11] and quantum lithography[12, 13], etc. However, by the Rayleigh’s criterion[14], when the widths of two closely located point sources are larger than the separation of the two sources, it becomes more difficult to distinguish between the two sources and the measurement precision of the separation vanishes using the classical direct imaging method. This imposes a severe limitation on the resolution of imaging.
The Rayleigh’s criterion indicates the insufficiency of direct measurement for high-resolution imaging in the positional space of optical sources. To precisely measure the centroid separation of two point sources, it is critical to find a suitable measurement basis beyond the positional basis and a proper estimation method so that the imaging precision can surpass the classical precision limit of direct imaging. The quantum parameter estimation theory provides the mathematical tool to solve this problem. The quantum parameter estimation theory is rooted in its classical counterpart. The classical parameter estimation theory optimizes the estimation precision of unknown parameters over all possible estimation strategies and gives the optimal estimation precision for parameters in a given parameter-dependent probability distribution by the well-known Cramér-Rao’s bound. The quantum parameter estimation theory considers the parameters in a quantum system, and further optimizes the estimation precision over all possible measurement basis in addition to the optimization of estimation strategies, since different quantum measurement bases lead to different probability distributions of the measurement results and thus to different estimation precisions of the unknown parameters. The precision limit of quantum estimation is characterized by the quantum Cramér-Rao bound, which is determined by the inverse of the quantum Fisher information matrix [15, 16, 17, 18, 19].
Inspired by the quantum estimation theory, Tsang et al. proposed a technique of superresolution for two incoherent optical sources [20] to break the Rayleigh’s curse. They found that by optimizing the measurement basis the precision of estimating the centroid separation between the two sources does not vanish when the separation approaches zero. This is in sharp contrast to the Rayleigh’s curse, which shows the power of quantum estimation theory and opens up a new avenue for precision resolution of two incoherent optical point sources.
@@
An implicit assumption in the original superresolution protocol is that the photon intensities of the two incoherent optical sources are known and equal. This assumption looks trivial at first glance, but turns out to play a critical role in the superresolution: if the photon numbers are unknown and need to be estimated simultaneously with the centroid separation of the two sources, the estimation precision of the centroid separation will vanish as the separation approaches zero if the photon intensities are not actually equal and the Rayleigh’s curse resurges [85]. As the photon intensities could vary or fluctuate due to noises and technical imperfections in reality, this imposes challenges to the application of superresolution in practice. In recent years, the role of partial coherence in the superresolution has also been investigated intensively [86, 87, 88, 89, 90, 75, 91, 92], and entanglement has been found useful to enhance the precision of superresolution [93, 94].
In this work, we analyze the effect of unknown photon numbers on the precision of superresolution from the viewpoint of quantum multiparameter estimation in detail, and obtain the constraint condition on the difference of the photon numbers between the two incoherent point sources to overcome the resurgence of Rayleigh’s curse. The results show the precision may deviate from that with identical photon numbers, depending on the magnitude of the difference of the photon numbers. We further consider two incoherent point sources with different point-spread functions, and find that the competition between the difference of the point-spread functions and the difference of the photon numbers determines the precision of superresolution for the centroid separation. We analyze the competition between these parameters in detail, and show that the superresolution can have various precision limits depending on the magnitudes of these parameters, beyond the precision limit of two optical sources with identical photon numbers and point-spread functions.
@@ II Preliminaries
Suppose we have two point sources imaged on a one-dimensional plane with unknown centroid positions $X_{1}$ and $X_{2}$. If the photon numbers of the two sources are $N_{1}$ and $N_{2}$ respectively, the state of the two sources can be represented by a quantum density operator [95]
@@
Here, $\epsilon$ is the ratio between the difference and the total of the photon numbers of the two optical sources,
$\epsilon=\frac{N_{2}-N_{1}}{N_{1}+N_{2}},$ (2)
$|\psi\rangle$ is a state defined by a point-spread function on the one-dimensional imaging plane assumed to be real and symmetric about the origin point [96] throughout this paper,
@@
and ${\rm e}^{-iX_{k}\hat{P}}$, $i=1,2$, are the displacement operators with $\hat{P}$ defined as
@@
Now, the question is: how precise can one measure and estimate the centroid separation,
@@ II.1 Direct imaging and Rayleigh’s curse
The most straightforward method to measure the centroid separation is directly detecting and counting photons arrived on the image plane from the two point sources [97]. The position distribution of photons on the image plane is
In order to quantify the precision limit of estimating the centroid separation of two incoherent optical sources, the well-known Cramér-Rao bound [15, 18, 19, 23] from the classical parameter estimation theory needs to be invoked, which is reviewed in Appendix A. For the task of resolving two point sources, the centroid positions of the two sources, $X_{1}$ and $X_{2}$, or equivalently the mean and difference of the two centroid positions, $\bar{X}=\frac{X_{1}+X_{2}}{2}$ and $d=X_{2}-X_{1}$ are unknown, so $\bar{X}$ and $d$ are the parameters to estimate.
@@
where $N_{{\rm tot}}$ is total number of photons from the two optical sources,
@@
For the direct imaging with an ideal continuum photodetector on the image plane, if the intensities and the point-spread functions of the two optical sources are the same, i.e., $N_{1}=N_{2}=N$, the position distribution of the arrival photons can be written as
@@
When the centroid separation $d$ is small, it can be shown by the Cramér-Rao bound 7 that the estimation precision of the centroid separation $d$ by direct imaging is
$\mathcal{H}_{d}^{({\rm direct})}=N_{{\rm tot}}/(\mathcal{J}^{-1})_{22}=\frac{1}{16}d^{2}\alpha_{2}N_{\text{tot}}.$ (12)
The detail of derivation is provided in Appendix A. Obviously, when $d\rightarrow 0$, the estimation separation precision vanishes, so one can hardly obtain any information about the centroid separation by direct imaging when the two point sources are sufficiently close to each other. This is the Rayleigh’s curse for direct imaging of two identical optical point sources.
The above result can be generalized to the case that the two point-spread functions are the same but the photon intensities are unequal. In this case, the position distribution of the photons can be written as
@@
The estimation precision of centroid separation by direct imaging can be found to be
@@
(See Appendix A for detail.) This indicates that one cannot get any information about the centroid separation $d$ by direct imaging in the limit $d\rightarrow 0$ in the generalized case either.
II.2 Quantum superresolution
To overcome the Rayleigh’s curse, one needs to further resort to the quantum estimation theory, which optimizes the Cramér-Rao bound over all possible quantum measurement schemes, including generalized measurements that are described by Kraus operators [98].
For the current quantum superresolution problem, the centroid positions of the two point sources, $X_{1}$ and $X_{2}$, or equivalently the mean and difference of the two centroid positions, $\bar{X}=\frac{X_{1}+X_{2}}{2}$, and $d=X_{2}-X_{1}$ are unknown, so $\bar{X}$ and $d$ are the parameters to estimate.
The quantum estimation theory for the quantum superresolution problem is briefly reviewed in Appendix B, which tells that the estimation precision of $\bar{X}$ and $d$ is lower bounded by the following quantum Cramér-Rao’s inequality,
@@ II.2 Quantum superresolution
by recognizing that the quantum Cramér-Rao bound is equivalent to the Holevo bound when this condition is satisfied and the latter is always achievable asymptotically with a sufficiently large number of states [99, 85, 100].
For the current problem of quantum superresolution, the interested parameter is the centroid separation $d$, so one can choose
@@
assuming $g_{1}=\bar{X}$ and $g_{2}=d$, and ${\rm Tr}(WC)$ gives the estimation variance of the centroid separation. And the above saturation condition can indeed be satisfied by the current problem, due to the reality of $\psi(x)$ which makes the symmetric logarithmic derivatives ^$\hat{\mathcal{L}}_{i}$ and ^$\hat{\mathcal{L}}_{j}$ also real according to Eq. (17) and thus the compatibility condition ${\rm ImTr}(\hat{\rho}\hat{\mathcal{L}}_{i}\hat{\mathcal{L}}_{j})=0$ can be satisfied. So the quantum Cramér-Rao bound provides an achievable precision limit for the estimation of centroid separation allowed by quantum mechanics.
A striking result found by Tsang et al. [20] is that the best estimation precision of the centroid separation $d$ over all quantum measurements for two incoherent point sources with the same photon numbers, determined by the quantum Cramér-Rao’s bound, keeps nonzero even when the separation goes to zero, implying the Rayleigh’s curse is no longer effective when the quantum measurement is optimized. The optimal precision of estimating the centroid separation $d$ in the limit $d\rightarrow 0$ is found to be
@@
An interesting finding in recent years is the resurgence of Rayleigh’s curse when the photon numbers $N_{1}$, $N_{2}$ of two point sources are unknown and need to be estimated as well [85]. In this case, one needs to estimate three parameters, $\bar{X},\,d,\,\epsilon$, simultaneously. If the photon numbers of the two point sources are different, i.e., $\epsilon\neq 0$, it was found that up to the second order of $d$, the precision of estimating $d$ was
@@
It can be immediately seen that in this case, $\mathcal{H}_{d}$ is of order $O(d^{2})$ and will vanish as $d$ goes to zero, implying the superresolution no longer works and the Rayleigh’s curse occurs again. This is the resurgence of Rayleigh’s curse for two incoherent point sources with unknown and unbalanced photon intensities [85, 101, 93].
@@ III Superresolution for two identical point sources
In this work, we investigate the resurgence of Rayleigh’s curse in detail and study the condition for the superresolution to avoid the the resurgence of Rayleigh’s curse.
As a finite difference between the photon numbers of the two sources can cause the resurgence of Rayleigh’s curse while the equality between the two photon numbers leads to superresolution, we are interested in the condition on the photon numbers to realize the superresolution: is it possible to realize the superresolution with a sufficiently small but nonzero difference between the photon numbers of the two sources and how close the two photon numbers need to be? In addition to the photon numbers, we will also investigate the effect of slight difference between the point-spread functions of the two optical sources on suppressing the resurgence of Rayleigh’s curse, and show its competition with the difference of the photon numbers in determining the precision of resolving the two point sources.
We consider two identical incoherent point sources located at $X_{1}$, $X_{2}$ with unknown photon numbers $N_{1}$, $N_{2}$ in this section. The unknown parameters to estimate in this case is $\bar{X}$, $d$, $\epsilon$, and $\epsilon$ is the ratio between the difference and the total of the photon numbers defined in Eq. (2). The density matrix $\hat{\rho}$ of the two sources is given by Eq. (1). By working out the eigenvalues and eigenstates of $\hat{\rho}$ and applying the formula of the quantum Fisher information matrix, the quantum Fisher information matrix with respect to these three unknown parameters can be obtained as
@@
According to the quantum Cramér-Rao bound, the estimation precision of the centroid separation $d$ can be obtained as
@@
To investigate the mechanism underlying the resurgence of Rayleigh’s curse when the photon numbers $N_{1}$ and $N_{2}$ of the two sources are different, we expand the numerator and denominator of the precision $\mathcal{H}_{d}$ (27) to the fourth order of $d$ respectively, and the result turns out to be
@@
Obviously, when the photon intensities of the two optical sources are unequal, i.e., $N_{1}\neq N_{2}$, the coefficient $D_{1}$ is finite, so $\mathcal{H}_{d}=O(d^{2})$, thus $\mathcal{H}_{d}$ vanishes when $d\rightarrow 0$, which is the resurgence of the Rayleigh’s curse. In contrast, when the photon intensities of the two optical sources are equal, i.e., $\epsilon\rightarrow 0$ and $N_{{\rm tot}}=N_{1}+N_{2}=2N$, $D_{1}=0$, so the estimation precision of $d$ (29) becomes
@@
Nevertheless, a further observation of $\mathcal{H}_{d}$ in Eq. (29) suggests that even when $N_{1}\neq N_{2}$ there is still opportunity to have nonvanishing estimation precision for the centroid separation $d$ if the coefficient $D_{1}$ is sufficiently small. It can be seen that if $D_{1}$ has the order $O\left(d^{2}{\rm Var}(\hat{P}^{2})\right)$, i.e., the order of $\epsilon$ is $O\left(d\sqrt{{\rm Var}(\hat{P}^{2})/\langle\hat{P}^{2}\rangle}\right)$, the precision $\mathcal{H}_{d}$ keeps finite when $d$ is small, implying the quantum superresolution scheme can still work in this case. Furthermore, if the order of the coefficient $D_{1}$ is higher than $O(d^{2}{\rm Var}(\hat{P}^{2}))$, or equivalently the order of $\epsilon$ is higher than $O\left(d\sqrt{{\rm Var}(\hat{P}^{2})/\langle\hat{P}^{2}\rangle}\right)$, $D_{1}$ can be dropped when $d\rightarrow 0$, and $\mathcal{H}_{d}$ is equal to $C/D_{2}=N_{{\rm tot}}\langle\hat{P}^{2}\rangle$ which restores the original result for the precision of superresolution in [20].
The above observation immediately leads to the condition for the quantum superresolution of two identical incoherent optical sources: the order of the ratio $\epsilon$ between the difference and the total of the photon numbers of the two optical sources must be higher than or equal to $O\left(d\sqrt{{\rm Var}(\hat{P}^{2})/\langle\hat{P}^{2}\rangle}\right)$. The original result in [20] was obtained for the case $N_{1}=N_{2}$, but is effective for the range of $\epsilon$ that is higher than $O\left(d\sqrt{{\rm Var}(\hat{P}^{2})/\langle\hat{P}^{2}\rangle}\right)$, while the superresolution scheme can actually work in a broader range including the case that the order of $D_{1}$ is equal to $O\left(d^{2}{\rm Var}(\hat{P}^{2})\right)$. But the latter has a lower precision than the former, as the ratio between the two precisions is

### v2

# Quantum-limited superresolution of two arbitrary incoherent point sources: beating the resurgence of Rayleigh’s curse

@@ Abstract
Superresolution has been demonstrated to overcome the limitation of the Rayleigh’s criterion and achieve significant improvement of the precision in resolving the separation of two incoherent optical point sources. However, in recent years, it was found that if the photon numbers of the two incoherent optical sources are unknown, the precision of superresolution vanishes when the two photon numbers are actually different. In this work, we first analyze the estimation precision of the separation between two incoherent optical sources with the same point-spread functions in detail, and show that when the photon numbers of the two optical sources are different but sufficiently close, the superresolution can still realized but with different precisions. We find the condition on how close the photon numbers of two optical sources need to be to realize the superresolution, and derive the precision of superresolution in different regimes of the photon number difference. We further consider the superresolution for two incoherent optical sources with different point-spread functions, and show that the competition between the difference of photon numbers, the difference of the two point-spread functions and the separation of the two optical sources determines the precision of superresolution. The results exhibit precision limits distinct from the case of two point sources with identical point-spread functions and equal photon numbers, and extend the realizable regimes of the quantum superresolution technique. The results are finally illustrated by Gaussian point-spread functions.
@@ I Introduction
Optical imaging is an important sensing technique that has contributed to various research areas, such as medical technology, astronomy, nanotechnology and biochemistry, etc. Typical imaging methods include classical and quantum imaging [1, 2], fluorescence imaging [3, 4, 5], the superlens [6, 7], two-photon microscopy [8, 9], quantum holography [10, 11] and quantum lithography[12, 13], etc. However, by the Rayleigh’s criterion[14], when the separation of the two sources is smaller than the widths of two closely located point sources, it becomes extremely difficult to distinguish between the two sources. This imposes a severe limitation on the resolution of imaging.
The Rayleigh’s criterion indicates the insufficiency of direct measurement for high-resolution imaging in the positional space of optical sources. To precisely measure the separation of two point sources, it is critical to find a suitable measurement basis beyond the positional basis and a proper estimation method so that the imaging precision can surpass the classical precision limit of direct imaging. The quantum parameter estimation theory provides the mathematical tool to solve this problem. The quantum parameter estimation theory is rooted in its classical counterpart. The classical parameter estimation theory optimizes the estimation precision of unknown parameters over all possible estimation strategies and gives the optimal estimation precision for parameters in a given parameter-dependent probability distribution by the well-known Cramér-Rao’s bound. The quantum parameter estimation theory considers the parameters in a quantum system, and further optimizes the estimation precision over all possible measurement basis in addition to the optimization of estimation strategies, since different quantum measurement bases lead to different probability distributions of the measurement results and thus to different estimation precisions of the unknown parameters. The precision limit of quantum estimation is characterized by the quantum Cramér-Rao bound, which is determined by the inverse of the quantum Fisher information matrix [15, 16, 17, 18, 19].
Inspired by the quantum estimation theory, Tsang et al. proposed a technique of superresolution for two incoherent optical sources [20] to surpass the Rayleigh’s criterion. While the measurement precision of the separation vanishes by the classical direct imaging method when the separation is close to zero, which is known as Rayleigh’s curse, the precision of estimating the separation of the two sources does not vanish by optimizing the measurement basis. This is in sharp contrast to the Rayleigh’s criterion, which shows the power of quantum estimation theory and opens up a new avenue for precision resolution of two incoherent optical point sources.
@@
An implicit assumption in the original superresolution protocol is that the intensities of the two incoherent optical sources are known and equal. This assumption looks trivial at first glance, but turns out to play a critical role in the superresolution: if the photon numbers are unknown and need to be estimated simultaneously with the separation of the two sources, the superresolution can work only when the intensities are equal and the estimation precision of the separation will vanish otherwise [85]. As the intensities could vary or fluctuate due to noises and technical imperfections in reality, this imposes challenges to the application of superresolution in practice. In recent years, the role of partial coherence in the superresolution has also been investigated intensively [86, 87, 88, 89, 90, 75, 91, 92], and entanglement has been found useful to enhance the precision of superresolution [93, 94].
In this work, we analyze the effect of unknown photon numbers on the precision of superresolution from the viewpoint of quantum multiparameter estimation in detail, and obtain the constraint condition on the difference of the photon numbers between the two incoherent point sources to avoid the vanishing of estimation precision. The results show the precision may deviate from that with identical photon numbers, depending on the magnitude of the difference of the photon numbers. We further consider two incoherent point sources with different point-spread functions, and find that the competition between the difference of the point-spread functions and the difference of the photon numbers determines the precision of superresolution for the separation. We analyze the competition between these parameters in detail, and show that the superresolution can have various precision limits depending on the magnitudes of these parameters, beyond the precision limit of two optical sources with identical photon numbers and point-spread functions.
@@ II Preliminaries
Suppose two incoherent optical pointlike sources are located at unknown positions $X_{1}$ and $X_{2}$ on the one-dimensional object plane. Neglecting the probability that more than one photon simultaneously arrives at the image plane in the Poisson limit, the imaging process can be regarded as the events of detecting a series of single photons over multiple coherence time intervals of the sources [95]. Assume the intensities of the two optical sources can be different. Denote the average numbers of photons emitted by the two sources within a given observation time as $N_{1}$ and $N_{2}$ respectively. The average single-photon state of the optical field arriving at the image plane can be described by a density operator $\hat{\rho}$ [96],
@@
The state $|\psi\rangle$ is defined as
@@
where $\psi(x)$ is the point-spread functions of the imaging system assumed to be real and symmetric about the origin point [97] throughout this paper, $\epsilon$ is the ratio between the difference and the total of the photon numbers of the two optical sources,
$\epsilon=\frac{N_{2}-N_{1}}{N_{2}+N_{1}},$ (3)
and ${\rm e}^{-iX_{i}\hat{P}}$, $i=1,2$, are the displacement operators with $\hat{P}$ defined as
@@
The probability distribution of a single photon arriving on the image plane is [20]
$\Lambda(x)=\frac{1-\epsilon}{2}\left|\langle x|{\rm e}^{-iX_{1}\hat{P}}|\psi\rangle\right|^{2}+\frac{1+\epsilon}{2}\left|\langle x|{\rm e}^{-iX_{2}\hat{P}}|\psi\rangle\right|^{2},$ (5)
and the photon count at a width $dx$ around the position $x$ of the detector is approximately $(N_{1}+N_{2})\Lambda(x)dx$.
Now, the question is: how precise can one measure and estimate the separation,
@@
The well-known Rayleigh’s criterion in optical imaging and the quantum superresolution scheme proposed in the seminal paper [20] concern the distinguishability of two pointlike optical sources with identical point-spread functions up to a small spatial translation, for which the quantum model is given by Eq. (1). In this paper, we will revisit the quantum superresolution scheme for two point sources with identical point-spread functions to investigate the validity condition of quantum superresolution, and extend it to the point sources with different point-spread functions in Sec. IV.
@@ II.1 Direct imaging and Rayleigh’s curse
The most straightforward method to measure the separation is directly detecting and counting photons arrived on the image plane from the two point sources [98].
In order to quantify the precision limit of estimating the separation of two incoherent optical sources, the well-known Cramér-Rao bound [15, 18, 19, 23] from the classical parameter estimation theory needs to be invoked, which is reviewed in Appendix A. For the task of resolving two point sources, the positions of the two sources, $X_{1}$ and $X_{2}$, or equivalently the centroid $\bar{X}=\frac{X_{1}+X_{2}}{2}$ and the separation $d=X_{2}-X_{1}$ of the two sources, are unknown, so $\bar{X}$ and $d$ are the parameters to measure and estimate.
@@
where $N_{{\rm tot}}$ is total number of photons from the two optical sources within a given detection time,
@@
For the direct imaging with an ideal continuum photodetector on the image plane, if the intensities and the point-spread functions of the two optical sources are the same, i.e., $N_{1}=N_{2}=N$, the probability distribution of a single photon arrived on the image plane can be written as
@@
When the separation $d$ is small, it can be shown [72] by the Cramér-Rao bound (7) that the estimation precision of the separation $d$ by direct imaging is
$\mathcal{H}_{d}^{({\rm direct})}=N_{{\rm tot}}/(\mathcal{J}^{-1})_{22}=\frac{1}{8}d^{2}\alpha_{2}N,$ (12)
where
[…]
The detail of derivation is provided in Appendix A. Obviously, when $d\rightarrow 0$, the estimation separation precision vanishes, so one can hardly obtain any information about the separation by direct imaging when the two point sources are sufficiently close to each other. This is the Rayleigh’s curse for direct imaging of two identical optical point sources [20].
The above result can be generalized to the case that the two point-spread functions are the same but the intensities are unequal. In this case, the probability distribution of a single photon on the image plane can be written as
@@
The estimation precision of the separation by direct imaging can be found [72] to be
@@
(See Appendix A for detail.) This indicates that one cannot get any information about the separation $d$ by direct imaging in the limit $d\rightarrow 0$ in the generalized case either.
II.2 Quantum-limited superresolution
To surpass the Rayleigh’s criterion, one needs to further resort to the quantum estimation theory, which optimizes the Cramér-Rao bound over all possible quantum measurement schemes, including generalized measurements that are described by Kraus operators [99].
For the current quantum superresolution problem, the positions of the two point sources, $X_{1}$ and $X_{2}$, or equivalently the centroid $\bar{X}=\frac{X_{1}+X_{2}}{2}$ and the separation $d=X_{2}-X_{1}$ are unknown, so $\bar{X}$ and $d$ are the parameters to estimate.
The quantum estimation theory for the superresolution problem is briefly reviewed in Appendix B, which tells that the estimation precision of $\bar{X}$ and $d$ is lower bounded by the following quantum Cramér-Rao’s inequality,
@@ II.2 Quantum-limited superresolution
by recognizing that the quantum Cramér-Rao bound is equivalent to the Holevo bound when this condition is satisfied and the latter is always achievable asymptotically with a sufficiently large number of states [100, 101].
For the current problem of quantum superresolution, the interested parameter is the separation $d$, so one can choose
@@
assuming $g_{1}=\bar{X}$ and $g_{2}=d$, and ${\rm Tr}(WC)$ gives the estimation variance of the separation. And the above saturation condition can indeed be satisfied by the current problem, due to the reality of $\psi(x)$ which makes the symmetric logarithmic derivatives ^$\hat{\mathcal{L}}_{i}$ and ^$\hat{\mathcal{L}}_{j}$ also real according to Eq. (18) and thus the compatibility condition ${\rm ImTr}(\hat{\rho}\hat{\mathcal{L}}_{i}\hat{\mathcal{L}}_{j})=0$ can be satisfied. So the quantum Cramér-Rao bound provides an achievable precision limit for the estimation of the separation allowed by quantum mechanics.
A striking result found by Tsang et al. [20] is that the best estimation precision of the separation $d$ over all quantum measurements for two incoherent point sources with the same photon numbers, determined by the quantum Cramér-Rao’s bound, keeps nonzero even when the separation goes to zero, implying the Rayleigh’s criterion is violated when the quantum measurement is optimized. The optimal precision of estimating the separation $d$ in the limit $d\rightarrow 0$ is found to be
@@
An interesting finding in recent years is the vanishing of estimation precision when the photon numbers $N_{1}$, $N_{2}$, of two point sources are unknown and need to be estimated as well [85]. In this case, one needs to estimate three parameters, $\bar{X},\,d,\,\epsilon$, simultaneously. If the photon numbers of the two point sources are different, i.e., $\epsilon\neq 0$, it was found that up to the second order of $d$, the precision of estimating $d$ was
@@
It can be immediately seen that in this case, $\mathcal{H}_{d}$ is of order $O(d^{2})$ and will vanish as $d$ goes to zero, implying the superresolution no longer works and the Rayleigh’s curse applies again. This is sometimes termed as the resurgence of Rayleigh’s curse for two incoherent point sources with unknown and unbalanced intensities [85, 102].
@@ III Superresolution for two identical point sources
In this work, we investigate the problem of precision vanishing in resolving two incoherent point sources with unknown and unequal intensities and study the condition for the quantum superresolution to address this problem in detail.
As a finite difference between the photon numbers of the two sources can cause the resurgence of Rayleigh’s curse while the equality between the two photon numbers leads to superresolution, we are interested in the condition on the photon numbers to realize the superresolution: is it possible to realize the superresolution with a sufficiently small but nonzero difference between the photon numbers of the two sources and how close the two photon numbers need to be? In addition to the photon numbers, we will also investigate the effect of slight difference between the point-spread functions of the two optical sources on suppressing the Rayleigh’s criterion, and show its competition with the difference of the photon numbers in determining the precision of resolving the two point sources.
We consider two identical incoherent point sources located at $X_{1}$, $X_{2}$ with unknown photon numbers $N_{1}$, $N_{2}$ in this section. The unknown parameters to estimate in this case is $\bar{X}$, $d$, $\epsilon$, and $\epsilon$ is the ratio between the difference and the total of the photon numbers defined in Eq. (3). The density matrix $\hat{\rho}$ of a single photon is given by Eq. (1). By working out the eigenvalues and eigenstates of $\hat{\rho}$ and applying the formula of the quantum Fisher information matrix, the quantum Fisher information matrix with respect to these three unknown parameters can be obtained [85] as
@@
According to the quantum Cramér-Rao bound, the estimation precision of the separation $d$ can be obtained as
@@
To investigate the mechanism underlying the vanishing of estimation precision when the photon numbers $N_{1}$ and $N_{2}$ of the two sources are different, we expand the numerator and denominator of the precision $\mathcal{H}_{d}$ (28) to the fourth order of $d$ respectively, and the result turns out to be
@@
Obviously, when the intensities of the two optical sources are unequal, i.e., $N_{1}\neq N_{2}$, the coefficient $D_{1}$ is finite, so $\mathcal{H}_{d}=O(d^{2})$, thus $\mathcal{H}_{d}$ vanishes when $d\rightarrow 0$, and the Rayleigh’s curse takes effect again. In contrast, when the intensities of the two optical sources are equal, i.e., $\epsilon\rightarrow 0$ and $N_{{\rm tot}}=N_{1}+N_{2}=2N$, $D_{1}=0$, so the estimation precision of $d$ (30) becomes
@@
Nevertheless, a further observation of $\mathcal{H}_{d}$ in Eq. (30) suggests that even when $N_{1}\neq N_{2}$ there is still opportunity to have nonvanishing estimation precision for the separation $d$ if the coefficient $D_{1}$ is sufficiently small. It can be seen that if $D_{1}$ has the order $O\left(d^{2}{\rm Var}(\hat{P}^{2})\right)$, i.e., the order of $\epsilon$ is $O\left(d\sqrt{{\rm Var}(\hat{P}^{2})/\langle\hat{P}{}^{2}\rangle}\right)$, the precision $\mathcal{H}_{d}$ keeps finite when $d$ is small, implying the superresolution scheme can still work in this case. Furthermore, if the order of the coefficient $D_{1}$ is higher than $O(d^{2}{\rm Var}(\hat{P}^{2}))$, or equivalently the order of $\epsilon$ is higher than $O\left(d\sqrt{{\rm Var}(\hat{P}^{2})/\langle\hat{P}{}^{2}\rangle}\right)$, $D_{1}$ can be dropped when $d\rightarrow 0$, and $\mathcal{H}_{d}$ is equal to $C/D_{2}=N_{{\rm tot}}\langle\hat{P}{}^{2}\rangle$ which restores the original result for the precision of superresolution in [20].
The above observation immediately leads to the condition for the superresolution of two identical incoherent optical sources: the order of the ratio $\epsilon$ between the difference and the total of the photon numbers of the two optical sources must be higher than or equal to $O\left(d\sqrt{{\rm Var}(\hat{P}^{2})/\langle\hat{P}{}^{2}\rangle}\right)$. The original result in [20] was obtained for the case $N_{1}=N_{2}$, but is effective for the range of $\epsilon$ that is higher than $O\left(d\sqrt{{\rm Var}(\hat{P}^{2})/\langle\hat{P}{}^{2}\rangle}\right)$, while the superresolution scheme can actually work in a broader range including the case that the order of $D_{1}$ is equal to $O\left(d^{2}{\rm Var}(\hat{P}^{2})\right)$. But the latter has a lower precision than the former, as the ratio between the two precisions is